OpenStudy (anonymous):

The sum of three consecutive integers is 252. Find the integers. Define the variables and set up an equation.

OpenStudy (jlastino):

let the first integer be x the next two consecutive integers would be (x+1) and (x+2) The sum of the three is 252 so x+x+1+x+2=252 3x+3=252 3x=252-3 3x=249 x=249/3 x=83 Since you have found the first integer...what are the two numbers that come after 83?

OpenStudy (anonymous):

x+x+1+x+2 =252 solve for x